713-2N and 3N Upgrade all bay lights with LED lamps.…
704-1N Install speakers and route cable.…
Repair Sheetrock Control Joint at 731-2N.…
This EEC addresses planned renovation activities within 704-L.…
Poles H3B545 and H3B546 have been reported bad by Osmose and must be replaced.…
This EEC will replace the terminal block that was found broken during troubleshooting at the 12W sump…
722-4A Universal Wiring upgrades.…
Install Keri Door (Number 1) Infrastructure at 702-A.…
Funding will support the project team's small-scale research of methods to efficiently extract rare earth elements and platinum group metals from m…